the three digit number 2m3 is added to the number 326 to give the three digit.number 5n9,which is divisable by 9. find m+n
Found 2 solutions by FrankM, ankor@dixie-net.com:
Answer by FrankM(1040)

The rule of 9's (for divisibility) states the digits of a number must add to 9 or a multiple of 9 for the number to be divisible by 9.
Therefore for 5n9, n must be 4. 5+4+9=18
If n = 4, then m = 2

Answer by ankor@dixie-net.com(22740)

the three digit number 2m3 is added to the number 326 to give the three digit.number 5n9, which is dividable by 9. find m+n
:
2 m 3
3 2 6
-----
5 n 9
therefore
m + 2 = n
the digits of 5n has to be divisible by 9, 54 is the only one, n=4
m + 2 = 4
m = 2
m + n = 6